Given that P(B|A)=0.78 and P(A)=0.41, what is P(B AND A)? Round to three decimal places.

Provide your answer below
Answer #1

Solution:

Given that,

P(B | A) = 0.78

P(A) = 0.41

Therefore, By conditional Probability

P(B | A) = P(A B ) / P(A)

P(A B ) = P(B | A) * P(A)

= 0.78 * 0.41

= 0.320

P(B AND A) = 0.320